A gas is contained in a 3.0L container at a temperature of 25C. The gas exerts a pressure of 16atm on the container. If pressure is kept constant, what is the final volume of the gas if the temperature of the container is increased to 200C?

1
Expert's answer
2022-02-15T16:20:09-0500

Since pressure is kept constant, the only variable that is manipulated is temperature. This means that we can use Charles's law in order to compare volume and temperature. Since volume and temperature are on opposite sides of the ideal gas law, they are directly proportional to one another. As one variable increases, the other will increase as well.


Charles's law is written as follows:


"\frac{V_1}{T_1}=\frac{V_2}{T_2}"

To use this law, we must first convert the temperatures to Kelvin.


"25°C+273=298K\\



200°C+273=473K"


Use these temperatures and the initial volume to solve for the final volume.

"\frac{3L}{298K}



=\frac{V_2}

{473K}



\\V

_2

=4.8L"